Corporate valuation Scampini Technologies is expected to generate $200 million in free cash flow next year, and FCF is expected to grow at a constant rate of 8% per year indefinitely. Scampini has no debt or preferred stock, and its WACC is 12%. If Scampini has 35 million shares of stock outstanding, what is the stock's value per share? Round your answer to two decimal places. Each share of common stock is worth $ , according to the corporate valuation model.
Answer #1

Total firm value = FCF1 / required rate - growth rate

Total firm value = 200 / 0.12 - 0.08

Total firm value = 200 / 0.04

Total firm value = 5,000

Value per share = 5,000 / 35

Value per share = $142.86